OpenStudy (anonymous):

evaluate the function for x=-3 f(x)=4x-5

OpenStudy (anonymous):

f(-3)= 4(-3) -5 = -12-5 = -17

OpenStudy (anonymous):

thanks @mathtutoring22 :0

OpenStudy (anonymous):

No problem :)